# Scattering theory of higher order topological phases

### v1

@@ 4.2 Systems with magnetic rotation symmetry
A natural application of scattering invariants arises in network models, which provide an efficient way to compute transport properties of large systems. This makes network models a widely used platform to study localization–delocalization transitions in quantum Hall systems [33, 34] and higher-order topological insulators, as done for example in Refs. [35, 6, 36, 37]. Topological invariants in network models, however, have an ambiguous topological classification due to the absence of a bulk Hamiltonian. In this section, we demonstrate how the scattering topological invariant resolves the ambiguous topological classification of HOTIs network models by considering a two-dimensional $C_{4}\mathcal{T}$-symmetric topological superconductor. We use the network model introduced in Ref. [38], where $\mathcal{P}^{2}=1$ and $(C_{4}\mathcal{T})^{4}=-1$, to construct an annulus geometry and, similar to tight-binding models, attach leads to the inside and outside of the disk, as shown in Fig. 6(a). We construct the scattering geometry and compute the reflection matrix using a network models package [39].
@@
Our theory for scattering invariants relies on constructing a symmetric scattering geometry and attaching symmetric leads to it. We further demonstrate the universality of the procedure by applying it to a three-dimensional $C_{4}\mathcal{T}$-symmetric HOTI [40]. We consider the model introduced in Ref. [40]:
@@
where $\sigma_{i}$ and $\tau_{i}$ are Pauli matrices acting on the spin and orbital degrees of freedom, $k_{i}$ are the components of the wave vector, $q_{i}$ and $p$ are strengths of the Hamiltonian terms that break additional symmetries, and $\mu$ drives the system through the phase transition. Because the model is three-dimensional, the dimensional reduction procedure results in a two-dimensional effective Hamiltonian for the scattering invariant. Additionally, because $C_{4}\mathcal{T}$ symmetry rotates the system around the $z$-axis, we choose to keep $k_{z}$ and $\phi$ as the momentum parameters of the effective Hamiltonian, and $x$ and $y$ as the spatial coordinates of the scattering geometry. Therefore, we construct a translationally invariant cylindrical geometry with leads attached to the inside and outside of the cylinder using Kwant [32] and Qsymm [41], as shown in Fig. 7(a).
@@
We observe that both in the two-dimensional and three-dimensional cases, the topological invariant is determined by the reflection matrix at $\phi=0$. This is similar to an earlier work on antiferromagnetic topological insulators [43]. Such behavior occurs whenever an antiunitary symmetry is multiplied by a fractional translation, which makes the square of the symmetry operator change between different high symmetry lines in the Brillouin zone. In our construction, the same principle applies to fractional rotations. It also allows us to conclude that the topological classification of these systems with magnetic rotation exponentiating to $+1$ would be identical, except for the relevant value of $\phi$ moving in the reduced Brillouin zone.
@@ 4.3 Axion insulator: 3D inversion
Finally, in this section, we demonstrate how to apply the scattering invariant to point-group symmetries, which do not have a preferred direction for the leads or boundary modes. As an example, we consider a three-dimensional axion insulator [44, 45, 46], which is a second-order HOTI protected by inversion symmetry. In a finite sample, the axion insulator has a chiral mode encircling the entire sample [47], as shown in Fig. 8(a). Inversion symmetry forbids the chiral mode from being contracted to a point, but it does not constrain its position on the surface.
@@
Here $\sigma_{i}$ and $\tau_{i}$ are Pauli matrices acting on the spin and orbital degrees of freedom, and $k_{i}$ are the components of the wave vector. The Hamiltonian parameter $\alpha$ is the spin-orbit coupling strength, $\bm{B}$ is a magnetic field, which we use to break time-reversal symmetry, and $\mu$ is the chemical potential. This model is trivial for $\mu<0$ and topological for $\mu>0$. Because the HOTI is protected by inversion symmetry, we need to construct an inversion-symmetric scattering geometry to apply the scattering invariant. Such a geometry is a translationally-invariant cylinder where $k_{z}$ and $\phi$ are the momentum parameters of the effective Hamiltonian, and $x$ and $y$ are the spatial coordinates of the scattering geometry.
@@
$V_{\mathcal{I}}(\phi)r(\phi,k_{z})Q_{\mathcal{I}}^{\dagger}(\phi)=r(\phi,-k_{z}),\quad V_{\mathcal{I}}^{2}(\phi)=Q_{\mathcal{I}}^{2}(\phi)=e^{-i\phi},$ (21)
a commutation relation for $k_{z}=0,\pi$ that follows from Eqs. (5a) and (21). The equivalent $H_{\textrm{eff}}$ (3) is two-dimensional, and it has a combination of a glide symmetry and a sublattice symmetry. This topological phase was analyzed in Ref. [48], and its topological invariant relies on block-diagonalizing the reflection matrix by projecting it on the eigenbasis of $V_{\mathcal{I}}(\phi)$ and $Q_{\mathcal{I}}(\phi)$ at $k_{z}=0$ and $k_{z}=\pi$. The scattering invariant is:
@@
where $r_{+}(\phi,k_{z})$ is the reflection matrix projected onto the subspace of $V_{\mathcal{I}}(\phi)$ and $Q_{\mathcal{I}}(\phi)$ with eigenvalue $e^{i\phi/2}$. Its graphical representation is shown in Fig. 8(b), and we identify the numerator and denominator of the fraction as the way of defining $\sqrt{\det r(-\pi,0)}$ and $\sqrt{\det r(-\pi,\pi)}$ without the sign ambiguity. Similar to Sec. 4.1, we apply the procedure of App. D to ensure that the integrands are smooth functions of $\phi$. By applying the invariant to the scattering matrix of our example system, we confirm that it switches at the phase transition, as shown in Fig. 8(c-d). Fig. 8(d) shows an extended range of $\mu$ where the invariant flips sign due to the Fermi surface that appears for such values of $\mu$.
Scattering invariants are a powerful tool to study localization–delocalization transitions in disordered systems because they establish a direct connection between topology and transport properties. Reference [6], for example, studied the effect of the bulk delocalization transition of disordered axion insulators using a network model, and claimed that an earlier analysis [49, 50] of topological phases protected by an average symmetry does not extend to HOTIs. Our scattering invariant, however, allows us to formulate a proof of the existence of the delocalized phase in a disordered axion insulator along the lines of Ref. [49]. We do this by considering a finite but large cylindrical sample with disorder in one half being an inversion image of the disorder in the other half, such that locally disorder minimally breaks inversion symmetry, but globally the system is inversion symmetric. Because the scattering invariant must change across the phase transition, a mode must perfectly transmit between the inner and outer radius of the cylinder as a phase transition is crossed. Additionally, because the system is three-dimensional, this implies that the perfectly transmitted mode is accompanied by a metallic phase. This argument establishes that the axion insulator must have a delocalized phase in the presence of disorder that is symmetric on average.
@@ 5 Conclusion and discussion
In all cases, the scattering invariant correctly identified the topological phase transitions regardless of the detailed geometry of the scattering region. The topological invariants that we used apply a reasoning previously identified for antiferromagnetic topological insulators [43] or for topological crystalline insulators [48]. Additionally, our approach allows us to directly factor out a pure rotation that commutes with the other symmetries, a step that requires special treatment in the Hamiltonian approach [40].
Our examples together with general arguments indicate that our approach applies to all the second order HOTIs. On the other hand, higher order HOTIs, in particular three-dimensional phases with zero-dimensional protected modes, do not seem to fit into our framework because it is possible to choose the cylinder geometry in such a way that the corner modes do not appear on its inner or outer boundary. We expect that the need for holes with a flux in the scattering geometry will extend to higher order HOTIs. More specifically, we conjecture that third order HOTIs need a scattering geometry with two holes and two fluxes $\phi_{i}$ threaded through them such that the dimensional reduction procedure still maps to a $d-1$-dimensional $H_{\textrm{eff}}(\phi_{1},\phi_{2},\bm{k}_{d-3})$. The finite size gap of the surface states at the boundaries then still goes to zero as a function of the fluxes introduced.
Finally, our work reveals a different perspective on higher order topological insulators than presented in the literature. We have found that a nontrivial reflection matrix always corresponds to the appearance of a zero energy mode in the original model’s spectrum, as exemplified in the spectral flow of Fig. 4. Because the topology of $r$ may not change when the geometry is varied smoothly, we may contract the inner radius of the cylinder into a point and extend the outer radius to infinity. This demonstrates that all second order HOTIs have protected defect modes appearing at flux lines, which is a local property that was previously overlooked. It is interesting to explore if this property can be used to characterize disordered higher order topological phases, like those studied in Ref. [8].
@@ Acknowledgements
All authors contributed to the initial idea, the final project scope, identifying the topological invariants, writing code, and writing the manuscript. A. A. oversaw the project.
@@
This research was supported by the Netherlands Organization for Scientific Research (NWO/OCW) as part of the Frontiers of Nanoscience program, a NWO VIDI grant 016.Vidi.189.180.
@@ B.2 Antiunitary antisymmetry
where $\sqrt{Q}_{\mathcal{P}}$ and $\sqrt{V}_{\mathcal{P}}$ denote the square root of the matrix. At time-reversal invariant momenta we define $S^{\prime}={\sqrt{Q_{\mathcal{P}}}}^{\ast}S{\sqrt{V_{\mathcal{P}}}}$ and find $S^{\prime}=S^{\prime\ast}$ for $\mathcal{P}^{2}=1$, and $S^{\prime}=\sigma_{y}S^{\prime\ast}\sigma_{y}$ for $\mathcal{P}^{2}=-1$. The $\sigma_{y}$ operator is required to satisfy $\sqrt{Q}_{\mathcal{P}}\sqrt{Q}_{\mathcal{P}}^{\ast}=\sqrt{V}_{\mathcal{P}}\sqrt{V}_{\mathcal{P}}^{\ast}=-1$. This constraint may be incompatible with other constraints if multiple symmetries are present.
@@ B.3 Unitary antisymmetry
For example, in the presence of chiral symmetry Eq. (41) establishes that $S$ is isomorphic to a Hermitian matrix $S^{\prime}$. To find the transformation that makes $S$ Hermitian, we first observe that $Q_{\mathcal{C}}V_{\mathcal{C}}=V_{\mathcal{C}}Q_{\mathcal{C}}=\mathcal{C}^{2}=1$ from Eq. (39). Then, we use $Q_{\mathcal{C}}^{\dagger}=V_{\mathcal{C}}=1$ in Eq. (41) and find $S^{\prime}=S^{\prime\dagger}$ for $S^{\prime}(\phi,\bm{k}_{\textrm{d-2}})=V_{\mathcal{C}}(\phi,\bm{k}_{\textrm{d-2}})S(\phi,\bm{k}_{\textrm{d-2}})$.

### v2

@@ 4.2 Systems with magnetic rotation symmetry
A natural application of scattering invariants arises in network models, which provide an efficient way to compute transport properties of large systems. This makes network models a widely used platform to study localization–delocalization transitions in quantum Hall systems [34, 35] and higher-order topological insulators, as done, for example, in Refs. [36, 6, 37, 38]. Topological invariants in network models, however, have an ambiguous topological classification due to the absence of a bulk Hamiltonian. In this section, we demonstrate how the scattering topological invariant resolves the ambiguous topological classification of HOTI network models by considering a two-dimensional $C_{4}\mathcal{T}$-symmetric topological superconductor. We use the network model introduced in Ref. [39], where $\mathcal{P}^{2}=1$ and $(C_{4}\mathcal{T})^{4}=-1$, to construct an annulus geometry and, similar to tight-binding models, attach a lead to both the inside and outside of the disk, as shown in Fig. 6(a). We construct the scattering geometry and compute the reflection matrix using a network models package [40].
@@
Our theory for scattering invariants relies on constructing a symmetric scattering geometry and attaching symmetric leads to both the outside and inside surfaces. We further demonstrate the universality of this procedure by applying it to a three-dimensional $C_{4}\mathcal{T}$-symmetric HOTI [41] with protected hinge modes. We consider the model introduced in Ref. [41]:
@@
where $\sigma_{i}$ and $\tau_{i}$ are Pauli matrices acting on the spin and orbital degrees of freedom, $k_{i}$ are the components of the wave vector, $q_{i}$ and $p$ are strengths of the Hamiltonian terms that break additional symmetries, and $\mu$ drives the system through the phase transition. Because the model is three-dimensional, the dimensional reduction procedure results in a two-dimensional effective Hamiltonian for the scattering invariant. Additionally, because $C_{4}\mathcal{T}$ symmetry rotates the system around the $z$-axis, we choose to keep $k_{z}$ and $\phi$ as the momentum parameters of the effective Hamiltonian, and $x$ and $y$ as the spatial coordinates of the scattering geometry. Therefore, we use a translationally invariant cylindrical geometry with leads attached to the inside and outside of the cylinder, such that a total of eight hinge modes propagate along the outer and inner surfaces, as shown in Fig. 7(a). We construct the scattering geometry and model using Kwant [33] and Qsymm [42].
@@
We observe that both in the two-dimensional and three-dimensional cases, the topological invariant is determined by the reflection matrix at $\phi=0$. This is similar to an earlier work on antiferromagnetic topological insulators [44]. Such behavior occurs whenever an antiunitary symmetry is multiplied by a fractional translation, which makes the square of the symmetry operator change between different high symmetry lines in the Brillouin zone. In our construction, the same principle applies to fractional rotations. Thanks to this observation, we also conclude that the topological classification of systems with magnetic rotation exponentiating to $+1$ is identical, except that the relevant value of $\phi$ shifts in the reduced Brillouin zone.
@@ 4.3 Axion insulator: 3D inversion
Finally, in this section, we demonstrate how to apply the scattering invariant to point-group symmetries, which lack a preferred direction for the leads or boundary modes. As an example, we consider a three-dimensional axion insulator [45, 46, 47], which is a second-order HOTI protected by inversion symmetry. A finite axion insulator sample hosts a chiral mode that encircles the entire sample [48], as shown in Fig. 8(a). Inversion symmetry forbids the chiral mode from being contracted to a point, but it does not constrain its position on the surface.
@@
Here $\sigma_{i}$ and $\tau_{i}$ are Pauli matrices acting on the spin and orbital degrees of freedom, and $k_{i}$ are the components of the wave vector. The Hamiltonian parameter $\alpha$ is the spin-orbit coupling strength, $\bm{B}$ is a magnetic field, which we use to break time-reversal symmetry, and $\mu$ is the chemical potential. This model is trivial for $\mu<0$ and topological for $\mu>0$. Because the HOTI is protected by inversion symmetry, we construct an inversion-symmetric scattering geometry for applying the scattering invariant. Such a geometry is a translationally-invariant cylinder where $k_{z}$ and $\phi$ are the momentum parameters of the effective Hamiltonian, and $x$ and $y$ are the spatial coordinates of the scattering geometry.
@@
$V_{\mathcal{I}}(\phi)r(\phi,k_{z})Q_{\mathcal{I}}^{\dagger}(\phi)=r(\phi,-k_{z}),\quad V_{\mathcal{I}}^{2}(\phi)=Q_{\mathcal{I}}^{2}(\phi)=e^{-i\phi}.$ (21)
This follows from Eqs. (5a) and (21), and becomes a commutation relation for $k_{z}=0,\pi$. The equivalent $H_{\textrm{eff}}$ (3) is two-dimensional, and it has a combination of a glide symmetry and a sublattice symmetry. This topological phase was analyzed in Ref. [49], and its topological invariant relies on block-diagonalizing the reflection matrix by projecting it onto the eigenbasis of $V_{\mathcal{I}}(\phi)$ and $Q_{\mathcal{I}}(\phi)$ at $k_{z}=0$ and $k_{z}=\pi$. The scattering invariant is:
@@
where $r_{+}(\phi,k_{z})$ is the reflection matrix projected onto the subspace of $V_{\mathcal{I}}(\phi)$ and $Q_{\mathcal{I}}(\phi)$ with eigenvalue $e^{i\phi/2}$. Its graphical representation is shown in Fig. 8(b), and we identify the numerator and denominator of the fraction as the way of defining $\sqrt{\det r(-\pi,0)}$ and $\sqrt{\det r(-\pi,\pi)}$ without the sign ambiguity. Similar to Sec. 4.1, we apply the procedure of App. D to ensure that the integrands are smooth functions of $\phi$. By applying the invariant to the scattering matrix of our example system, we confirm that the invariant switches at the phase transition, as shown in Fig. 8(c-d). Fig. 8(d) shows an extended range of $\mu$ where the invariant flips sign due to the Fermi surface that appears for such values of $\mu$.
Scattering invariants are a powerful tool to study localization–delocalization transitions in disordered systems because they establish a direct connection between topology and transport properties. Reference [6], for example, studied the effect of the bulk delocalization transition of disordered axion insulators using a network model, and claimed that an earlier analysis [50, 51] of topological phases protected by an average symmetry does not extend to HOTIs. Our scattering invariant, however, allows us to formulate a proof of the existence of the delocalized phase in a disordered axion insulator along the lines of Ref. [50]. We do this by considering a finite but large cylindrical sample with disorder in one half being an inversion image of the disorder in the other half, such that locally disorder breaks inversion symmetry, but globally the system is inversion symmetric. Because the scattering invariant must change across the phase transition, a mode must perfectly transmit between the inner and outer radius of the cylinder as a phase transition is crossed. Additionally, because the system is three-dimensional, this implies that the perfectly transmitted mode is accompanied by a metallic phase. This argument establishes that the axion insulator must have a delocalized phase in the presence of disorder that is symmetric on average.
@@ 5 Conclusion and discussion
In all cases, the scattering invariant correctly identified the topological phase transitions regardless of the detailed geometry of the scattering region. The topological invariants that we used employ reasoning previously identified for antiferromagnetic topological insulators [44] or for topological crystalline insulators [49]. Additionally, our approach enables us to factor out a pure rotation that commutes with the other symmetries, while the Hamiltonian approach [41] requires special treatment of such symmetries.
Our examples, together with general arguments, indicate that our approach applies to all the second order HOTIs. On the other hand, higher order HOTIs, in particular three-dimensional phases with zero-dimensional protected modes, do not seem to fit into our framework. This is because one can choose the cylinder geometry so that the corner modes do not appear on its inner or outer boundary. We expect that the need for holes with a flux in the scattering geometry extends to higher order HOTIs. More specifically, we conjecture that third order HOTIs need a scattering geometry with two holes and two fluxes $\phi_{i}$ threaded through them, in such a way that the dimensional reduction procedure maps to a $d-1$-dimensional $H_{\textrm{eff}}(\phi_{1},\phi_{2},\bm{k}_{d-3})$. The finite size gap of the surface states at the boundaries then goes to zero as a function of the fluxes introduced.
Finally, our work reveals a different perspective on higher order topological insulators from those presented in the literature. We have found that a nontrivial reflection matrix always corresponds to the appearance of a zero energy mode in the original model’s spectrum, as exemplified in the spectral flow of Fig. 4. Because the topology of $r$ may not change when the geometry is varied smoothly, we may contract the inner radius of the cylinder into a point and extend the outer radius to infinity. This demonstrates that all second order HOTIs have protected defect modes appearing at flux lines, which is a local property that was previously overlooked. It is interesting to explore whether this property characterizes disordered higher order topological phases, like those studied in Ref. [8]. It is also interesting to apply our framework to identify quantized transport phenomena in HOTIs, likely related to the interplay between adiabatic pumping [52, 53] and symmetry eigenvalues.
@@ Acknowledgements
All authors contributed to the initial idea, final project scope, identification of the topological invariants, writing code, and writing the manuscript. A. A. oversaw the project.
@@
This research was supported by the Netherlands Organization for Scientific Research (NWO/OCW) as part of the Frontiers of Nanoscience program, and an NWO VIDI grant 016.Vidi.189.180.
@@ B.2 Antiunitary antisymmetry
where $\sqrt{Q}_{\mathcal{P}}$ and $\sqrt{V}_{\mathcal{P}}$ denote the square root of the matrix. At time-reversal invariant momenta we define $S^{\prime}={\sqrt{Q_{\mathcal{P}}}}^{\ast}S{\sqrt{V^{T}_{\mathcal{P}}}}$ and find $S^{\prime}=S^{\prime\ast}$ for $\mathcal{P}^{2}=1$, and $S^{\prime}=\sigma_{y}S^{\prime\ast}\sigma_{y}$ for $\mathcal{P}^{2}=-1$. The $\sigma_{y}$ operator is required to satisfy $\sqrt{Q}_{\mathcal{P}}\sqrt{Q}_{\mathcal{P}}^{\ast}=\sqrt{V}_{\mathcal{P}}\sqrt{V}_{\mathcal{P}}^{\ast}=-1$. This constraint may be incompatible with other constraints if multiple symmetries are present.
@@ B.3 Unitary antisymmetry
For example, in the presence of chiral symmetry Eq. (41) establishes that $S$ is isomorphic to a Hermitian matrix $S^{\prime}$. To find the transformation that makes $S$ Hermitian, we first observe that $Q_{\mathcal{C}}V_{\mathcal{C}}=V_{\mathcal{C}}Q_{\mathcal{C}}=\mathcal{C}^{2}=1$ from Eq. (39). Then, we use $Q_{\mathcal{C}}^{\dagger}=V_{\mathcal{C}}$ in Eq. (41) and find $S^{\prime}=S^{\prime\dagger}$ for $S^{\prime}(\phi,\bm{k}_{\textrm{d-2}})=V^{\dagger}_{\mathcal{C}}(\phi,\bm{k}_{\textrm{d-2}})S(\phi,\bm{k}_{\textrm{d-2}})$.